### v2

# Emergence of Light Cones in Long-range Interacting Spin Chains due to Destructive Interference

Abstract
We present a mechanism in long-range interacting Heisenberg spin chains at low temperatures that leads to the emergence of effective entanglement light cones. This mechanism arises from destructive interference among quantum effects that contribute to the entanglement of spins outside an identified light cone. As a result, entanglement remains suppressed in this region, facilitating the formation of effective entanglement light cones. Our analysis reveals that truncating the range of interactions weakens the destructive interference, resulting in an unexpected acceleration of entanglement transport along the chain. This prediction is proposed to be experimentally observable. Our work provides a fresh perspective on the physical mechanisms driving the emergence of light cones in long-range interacting quantum systems.
preprint: APS/123-QED
I Introduction
Understanding the propagation of entanglement in coherent, out-of-equilibrium many-body quantum systems is critical for both theoretical research and practical applications, especially in the context of recent advances linking entanglement to thermodynamics in closed quantum systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. In relativistic quantum field theory, the spread of correlations is strictly limited by a light cone [13]. However, in non-relativistic quantum systems, tunneling effects lead to a rapid build-up of correlations.
Despite the later circumstance, the Lieb-Robinson (LR) bound establishes an effective light cone structure in short-range interacting quantum systems, where correlations decay exponentially beyond this boundary [14]. The word effective emphasizes the existence of tunneling effects beyond this boundary [15]. Over the past few decades, the LR theorem has been refined [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] and validated [28, 29, 30, 31, 32] across a wide range of quantum systems, include long-range interacting systems; in this class of systems, due to the long-range nature of interactions, correlations can propagate arbitrarily fast, potentially leading to the absence of any light cones. However, various forms of light cones continue to persist in the long-range interacting regime [33, 22, 26, 34, 35, 36, 27, 37, 38, 39, 40, 26, 27].
In this work, we propose a new framework to explain the emergence of light cones in long-range interacting spin chains through destructive interference. Our theoretical analysis shows that destructive interference among quantum effects in Heisenberg spin chains at low temperatures leads to the formation of entanglement light cones. Specifically, while entanglement can propagate over long distances, interference cancels out the entangling effects outside an identified light cone, effectively suppressing entanglement in this region.
This mechanism is analogous to the destructive interference observed in quantum electrodynamics, where photons exceeding the speed of light undergo destructive interference, thus emergence of causal light cones at long distances [41].
Our theoretical and numerical analyzes demonstrate that when interactions are truncated, i.e., interactions beyond a certain cutoff length are omitted, the entanglement edge propagates faster compared to scenarios with full long-range interactions. This counter-intuitive effect is due to the weakening of the destructive interference in the case of truncated interactions. This prediction is proposed to be experimentally testable using quantum simulators such as trapped ions [42, 30, 43, 44, 45, 46].
Our results offer a new perspective on how quantum information propagates in long-range interacting systems, and why light cones persist to emerge in these systems. To the best of our knowledge, this novel framework, which attributes the emergence of light cones to the physical mechanism of destructive interference, has not been previously observed in the literature. This perspective enables a deeper exploration of LR bounds in long-range interacting quantum systems, and allows for further implications, useful to the quantum information processing community [27].
II Preliminary definitions and spin chain Model
As the primary model to explain our results, we examine the evolution of bipartite entanglement in Heisenberg spin-$\frac{1}{2}$ chains at low temperatures. This model is represented via bosons through the Holstein–Primakoff transformation [47]. At low temperatures, the density of excitations is low, making only the quadratic bosonic terms in the Hamiltonian relevant while higher-order terms can be ignored [48]. This approximation establishes the free-particle regime.
The Heisenberg spin $\frac{1}{2}$ chain Hamiltonian is given by
$\mathbf{H}=\frac{-1}{2}\sum_{j<l}^{N}\mathbf{J}(|j-l|)\mathbf{S}_{j}\mathbf{S}_{l}.$ (1)
Here $\mathbf{S}_{j}$ is the spin operator at site $j$. The interaction function $\mathbf{J}(n)$ is assumed to be real, symmetric $\mathbf{J}(n)=\mathbf{J}(-n)$, and absolutely convergent, i.e., $\sum_{n}|\mathbf{J}(n)|<\infty$. The special case of power-law decay interaction function is given by
$\mathbf{J}(n)=\mathbf{J}(|j-l|)=\frac{{J}}{|j-l|^{p}},$ (2)
with $J$ being the interaction amplitude and $p$ determining the interaction length. For simplicity, we set $\mathbf{J}(1)=1$, therefore $J=1$. We also set $\hbar=1$.
Initially, the system is in a ferromagnetic order. At $t=0$, we excite a low-density of spins, quenching the system’s state. These excitations travel throughout the chain as quasi-particles and lead to entanglement between subsystems. We measure bipartite entanglement between two continuous segments of the chain, labeled subsystems $M$ and $R$, to observe how the entanglement evolves. To measure bipartite entanglement, we use the linear bipartite entropy of subsystem $\mathcal{M}$ [49, 50], and denote this quantity by $\mathcal{Q}_{\mathcal{M}}(t)$.
Using the Taylor series expansion of the bipartite entanglement measure $\mathcal{Q}_{\mathcal{M}}(t)$ around $t=0$, we can write
$Q_{\mathcal{M}}(t)=\sum_{r=0}{\alpha_{r}}t^{2r}.$ (3)
The series only involves even exponents of $t$, because in the described setting we have $Q_{\mathcal{M}}(-t)=Q_{\mathcal{M}}(t)$ [51].
We can fully characterize entanglement through the coefficients $\alpha_{r}$. Let us denote by $q$ the distance between a local cause of entanglement, namely a local excitation of the system’s state, and where the effect, i.e., bipartite entanglement of $\mathcal{M}$, is measured.
In spin chains with Nearest-Neighbor (NN) interactions, we have $\alpha_{r}=0$ for $r<q$, and the first non-zero terms in the Taylor series expansion (3) is [51]
$\mathcal{Q}_{\mathcal{M}}(t)\propto(\frac{t}{q})^{2q}.$ (4)
This term predominates the entanglement expansion for $t<q$ [52]. (Note that since we set $\hbar=1$, and normalized the interaction function by $\mathbf{J}(1)=1$, then time $t$ and distance $q$ can be directly compared, such as in $t<q$). Also, entanglement $Q_{\mathcal{M}}(t)$ is exponentially suppressed for times before $t=q$, and grows after this time. Hence, equation (4), which corresponds to $r=q$ in expansion (3), governs early (here, $t\leq q$) entanglement growth at distance $q$.
As a result of this interconnection between $r$, $q$ and $t$, we can identify the entanglement light cone by the term $r=q$ in the Taylor series (3). Similarly, $r<q$ ($r>q$) denote to the outside (inside) of the light cone [53].
Analogously, when the coefficients $\alpha_{r}$ are vanishing (or exponentially suppressed) for $r<\frac{q}{2}$, and the early entanglement ($t<\frac{q}{2}$) follows
$\mathcal{Q}_{\mathcal{M}}(t)\propto(2t/q)^{q},$ (5)
we can identify the emerging light cone with $2r\approx q$ Also, $2r<q$ ($2r>q$) identifies the outside (inside) of this light cone.
In the next Sections, we will demonstrate the existence of the destructive interference mechanism outside the light cone $2r<q$.
III A Summarized mathematical demonstration of the destructive interference mechanism
Here, we will present a summary of the mathematical demonstration of the destructive interference phenomenon. The full proof is given in the Supplemental material [54]. This derivation demonstrates the underlying mechanism by which the entangling effects, contributing to the coefficients $\alpha_{r}$ in (3), destructively interfere and cancel out. Therefore, we show how $\alpha_{r}$ remains suppressed outside the light cone $2r<q$, despite the long-range nature of the interactions.
The parameters $\alpha_{r}$ in the Taylor series (3) are constructed by a linear combination of terms of the form
$\alpha_{r}\equiv\frac{1}{(2r)!}\sum_{r_{1}+r_{2}=2r}\binom{2r}{r_{1}}(-1)^{r_{1}}\mathbf{J}^{*r_{1}}(q)\mathbf{J}^{*r_{2}}(q),$ (6)
where $\mathbf{J}^{*r_{1}}(q)$ denotes $r_{1}$ consecutive convolutions of the interaction function $\mathbf{J}(n)$, evaluated at $q$. Therefore, showing that the terms on the r.h.s of (6) are vanishingly small proves that $\alpha_{r}\approx 0$ (in the appropriate scale described below). Here we focus on the case where subsystem $\mathcal{M}$ includes a spin at distance $q$ of the initial local quench.
Using the definition of the convolution operation we have:
$\begin{split}{{\mathbf{J}}^{*r}(q)}=&\sum_{\begin{subarray}{c}d_{1},d_{2},\cdots,d_{r}\in\mathbb{Z}-\{0\},\sum d_{i}=q\end{subarray}}\prod_{i}\mathbf{J}(d_{i}).\end{split}$ (7)
The condition $\sum d_{i}=q$ ensures that the summation variables $di$ sum to the total distance $q$, contributing to the expansion. All such combinations $\{d_{1},\cdots,d_{r}\}$ contribute to the expansion in (7). Here, we identify two general types of combinations $\{d_{1},\cdots,d_{r}\}$.
  • Type $1$: Most $d_{i}$’s are on the order of $1$ and only a few $d_{i}$’s are in the order of $r$.
  • Type $2$ includes the more uniform distributions of $d_{i}$’s, wherein more (than $1$) of the $d_{i}$’s are on the order of $r$.
For a wide classes of interaction functions, identified in [54], the Type $1$ combinations dominate the expansion (3), and the Type $2$ combinations are exponentially suppressed. This class includes power-law interactions with $p>2$. For these interactions, the expansion (7) can be written as:
$\begin{split}{{\mathbf{J}}^{*r}(q)}=&\frac{\binom{r}{1}}{(q-r+1)^{p}}+\frac{\binom{r}{2}}{\big{(}2(q-r)\big{)}^{p}}+\cdots.\end{split}$ (8)
(Here, the first term corresponds to $d_{1}=d_{2}=\cdots=d_{r-1}=1$ and $d_{r}=q-r+1$ (Type $1$). Since any of the $d_{i}$’s can be picked to equal $q-r+1$, there are $\binom{r}{1}$ such terms, which gives the numerator of the fraction. The second term in (8) corresponds to $d_{1}=d_{2}\cdots=d_{r-2}=1$, $d_{r-1}=2$, and $d_{r}=q-r$ (Type $1$). There exists $\binom{r}{2}$ of such terms).
Let us consider the first term in the expansion (8). A similar set of derivations holds for any other term in this series, so long as they correspond to a Type $1$ combination of $d_{i}$’s. When $0<r<q$, the first term in (8) can be written as:
$\begin{split}\frac{\binom{r}{1}}{(q-r+1)^{p}}=&\frac{r}{q^{p}}(1-\frac{r-1}{q})^{-p}\\ =&\frac{r}{q^{p}}\Big{(}1+p\frac{r-1}{q}+p(p+1)(\frac{r-1}{q})^{2}+\cdots\Big{)}.\end{split}$ (9)
If the first two terms in this Taylor expression are used as approximations for the high-order convolutions $\mathbf{J}^{*r_{1}}(q)$ and $\mathbf{J}^{*r_{2}}(q)$ in (6), then $\alpha_{r}$ can be expressed as (for $0<r_{1}<q$ and $0<r_{2}<q$)
$\begin{split}\alpha_{r}\equiv\frac{1}{(2r)!q^{2p}}\sum_{r_{1}+r_{2}=2r}\binom{2r}{r_{1}}(-1)^{r_{1}}&\Big{(}r_{1}+\frac{pr_{1}(r_{1}-1)}{q}\Big{)}\\ &\Big{(}r_{2}+\frac{pr_{2}(r_{2}-1)}{q}\Big{)}=0,\end{split}$ (10)
due to the binomial theorem. This equation manifests the destructive interference phenomenon. It demonstrates the cancellation between entangling effects at distance $q$, that leads to $\alpha_{r}=0$ up to the considered order of terms. Such cancellations occur for higher-order terms in both expansions (8) and (9), as well as for arbitrarily large subsystem $\mathcal{M}$ as shown in the full proof [54]; For higher-order terms in the expansion (9), the cancellation is not complete [55]. This leads to a marginal contribution to $\alpha_{r}$ from higher-order terms, which are characterized in the full proof [54]. Type $2$ combinations are exponentially suppressed and increasingly negligible for $p>2$. Also, note that since (10) vanishes independent of $p$, early entanglement dynamics is increasingly independent of the parameter $p$ when $p>2$. This effect is evident in our numerical simulation in [54].
In this summarized proof, we showed that the main body of entangling effects, travelling to distance $q$, destructively interfere and cancel out when $r_{1}<q$ and $r_{2}<q$. To ensure these inequalities, we can consider the region $2r<q$. Therefore, for the $2r<q$ light cone, by considering all higher-order terms in the expansions (8) and (9), the following theorem can be proved.
III.1 A theorem for $2r<q$ light cone
In the context described above, the following statement holds [54]:
Theorem: When $p>2$, outside the light cone $2r<q$ (with $q\gg 1$), we have $\alpha_{r}\sim{O}\big{(}\frac{p^{4}}{q^{2p+4}(2r-2)!}\big{)}$.
Let us elaborate on how this scaling is highly suppressed. Consider the case of NN interactions, the first non-zero $\alpha_{r}$ corresponds to $r=q$, and for this value we have $\alpha_{r}=\frac{1}{(r!)^{2}}$ [51]. In this case, the $r=q$ light cone is formed. Therefore, for the $2r\approx q$ light cone to form, the first non-zero $\alpha_{r}$ term (here $r\approx\frac{q}{2}$), must be of order $\hat{\alpha}_{r}=\frac{1}{(q/2)!^{2}}$. However, for $2r\approx q$, the scaling for $\alpha_{r}$ follows:
$\alpha_{r}\sim{O}\big{(}\frac{p^{4}}{q^{2p+4}(q-2)!}\big{)},$ (11)
which is significantly smaller than $\hat{\alpha}_{r}=\frac{1}{(q/2)!^{2}}$, since $\frac{\alpha_{r}}{\hat{\alpha}_{r}}=\frac{p^{4}}{q^{2p+4}}\frac{(q/2)!^{2}}{(q-2)!}\rightarrow 0$ for $q\gg 1$. (Note that $(q-2)!\gg(\frac{q}{2}!)^{2}$).
This theorem shows that for $p>2$, any possible light cone must be contained inside the light cone $2r>q$, as a result of the destructive interference between entangling effects.
IV Faster entanglement propagation for truncated interactions
The destructive interference between quasi-particles has a counter-intuitive implication: The entanglement edge must propagate faster when the interaction function is truncated, namely when the interaction terms that extend beyond the range ($\eta$) are omitted from the interaction function. This phenomenon is the direct consequence of quasi-particle cancellation in the early entanglement region due to the following reason. The destructive interference occurs when all entangling effects are superposed (see (10)). Therefore, if part of the stream of quasi-particles is obstructed due to truncation of the interaction function, then the destructive interference is weakened; Hence, the entanglement edge propagates faster. This phenomenon is mathematically proved, and numerically demonstrated in [54]. Here, a truncated interaction refers to:
$\mathbf{J}_{\eta}(n)=\begin{cases}\frac{J}{|n|^{p}},&n\in\{-\eta,\cdots,\eta\}\\ 0,&\text{otherwise},\end{cases}$ (12)
where $\eta=1$ gives nearest-neighbour interaction. We numerically study truncated interactions in [54], and confirm that destructive interference occurs when $p>2$.
As mentioned in the previous section, the structure of the emergent light cones are expected to be increasingly independent of $p$. This is due to the fact that (10) is vanishing, independent of $p$. In this regard, we numerically show [54] that the destructive interference phenomenon leads to the independence of the velocity of the entanglement propagation with respect to the exponent $p$. Based on this feature, and given that $p\rightarrow\infty$ corresponds to NN interactions, for $p>3$ light cones must be strictly linear (similar to the case of NN interactions), confirming the results in [22].
V Concluding remarks
In this work, we offer a new perspective on the well-studied problem of light cones in spin chains with long-range interactions, revealing new theoretical insights. We identify an inherent insulating behavior that limits entanglement propagation beyond the light cone. Specifically, we demonstrate that although the entangling effects propagate rapidly due to the long-range nature of the interactions, they are significantly suppressed by a destructive interference mechanism. While our primary analysis focuses on the long-range interacting Heisenberg chain, our findings also extend to free-particle spin chains. Additionally, in [54], we derive a necessary condition for general interaction functions, $\mathbf{J}(n)$—beyond the typical power-law forms—under which the destructive interference mechanism is active. We further show that truncating the interaction range weakens the destructive interference, resulting in faster entanglement propagation along the spin chain.
Our results show that the emergence of LR bounds in power-law decaying interactions with $p>2$ [56] coincides with the breakdown of destructive interference. Similarly, the presence of a strictly linear light cone for $p>3$ [22, 38] aligns with our finding that in this regime, destructive interference causes the entanglement propagation velocity to converge to that of NN interactions. The strong consistency of our results with existing research emphasizes the critical role that destructive interference plays in explaining the emergence of light cones.
We expect that the experimental observation of destructive interference among particles can be achieved through the use of truncated interactions. This prediction suggests an anomalous acceleration of entanglement growth in the presence of truncated interaction functions. We propose further exploration of experimental observation of this counter-intuitive prediction in several platforms that simulate many-body quantum systems, including cold atoms [28, 6] and trapped ions [30, 57, 45, 46]. Further analytical research should investigate the effects of destructive interference in higher-dimensional (two- and three-dimensional) lattices and assess the robustness of this phenomenon when higher-order bosonic operators are present. At higher temperatures, where quasi-particle interactions become significant, bound states may form [58, 30]. These bound states could alter the characteristics of entanglement propagation by introducing new features in the dynamics. Therefore, an important direction is to study the robustness of destructive interference in the presence of quasi-particle interactions.
Acknowledgements.
P.A would like to thank Hans Halvorson, Nicolo Defenu and Alessio Lerose for fruitful discussions. P.A acknowledges partial support from the Princeton Program in Plasma Science and Technology (PPST). H.R and P.A acknowledge partial support from the U.S Department Of Energy (DOE) grant (DE-FG02-02ER15344).